OpenStudy (anonymous):

How do you integrate x^2(1+2x^3)^5dx (borns are 1 to 0)

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^2(1+2x^3) dx ?\]

OpenStudy (anonymous):

i already figured it out but thanks!

OpenStudy (anonymous):

what about e^(x^1/2) / x^(1/2) borns are 4 to 1

OpenStudy (anonymous):

good, i left out the power of 5 anyways...

OpenStudy (anonymous):

let u = x^(1/2), so du = 1/(2x^(1/2))... you already have the factor of 1/(x^1/2)... got it?

OpenStudy (anonymous):

du = 1/(2x^(1/2)) dx*

OpenStudy (anonymous):

yes! thanks